2 . Problems 21-11 Assume that the reserve requirement is 12.5 percent. Also assume that banks do not hold excess reserves and there is no cash held by the public. The Federal Reserve decides that it wants to contract the money supply by $80 million using open-market operations.
Answer #1

Multiplier = 1/R

=1/0.125 = 8

The Fed wants to decrease the money supply by $80 million using the open market operations

So, it will sell 80/8 = $10 million worth of bonds.